OpenStudy (anonymous):

Find the seventh term of a sequence whose first term is 1 and whose common ratio is 3

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm a_1 \times r^{n-1}=a_n }\) in a geomatric sequence. In your case, \(\large\color{blue}{ \rm 1 \times 3^{(7-1)}=~? }\)

OpenStudy (solomonzelman):

geom ` e ` tric -:(